OpenStudy (anonymous):

PLEASE HELP! I understand the FOIL method but I don't get this. Write the product as a trinomial. (k – 1)(k – 4)

OpenStudy (anonymous):

What happens when you use the foil method?

OpenStudy (anonymous):

k^2-5k+4

OpenStudy (anonymous):

i believe, you just foil out and pretend the variables are constant.

OpenStudy (anonymous):

If you understand the foil method, then can you do it?

OpenStudy (anonymous):

Actually, my question is, do I keep the minus signs? and subtract it all once I'm done with the FOIL method

OpenStudy (anonymous):

Okay, so foil method gives you:\[ (k – 1)(k – 4) = k\times k+-1 \times k +k\times -4 + -1\times -4 \]

OpenStudy (anonymous):

Then you just have to simplify.

OpenStudy (anonymous):

\[ =k^2-k-4k+4 = k^2-5k+4 \]

OpenStudy (anonymous):

Refresh